OpenStudy (anonymous):

Please help! Hi, I'm stuck on a part of this question, we have to solve using Gauss Jordan Elimination: A corporation wants to lease a fleet of 12 airplanes with a combined carrying capacity of 220 passengers. The three available types of planes carry 10,15,20 passengers. How many of each type of plane should be leased?

OpenStudy (anonymous):

see assume like x,y,z 10x+15y+20z=220 x+y+z=12 and one more?

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

I understood these parts until I solved using Gauss Jordan Elimination and got: [1 0 1 | -8 ] [0 1 2 | 20 ] Here is the part where I didn't understand, x + z = -8 y + z = 20 z = r 8≤ z ≤10?

OpenStudy (anonymous):

oh ok.........wat did u not understand?

OpenStudy (anonymous):

The 8 ≤ z ≤ 10, I'm not sure if this should be given or we have to find it out.

OpenStudy (anonymous):

wat i thot was like we kno that x,y,z>=0 so....i thot we cn make smthin out of it and is that answer correct? i dont see how x+z=-8? it's like the company is giving away planes?

hero (hero):

Is there more than one solution for this?

OpenStudy (anonymous):

it's obvious doncha think

OpenStudy (anonymous):

They're leasing planes, and we need 12 planes, they should be a combination of 3 planes that carry either 10, 15 or 20 passengers. According to the answer there are multiple solutionsy, because it should be a combination. The final answer is; Either, z=8, x=0, y=4 or z=9, x=1, y=2, or z=10 x=2, y=0. The total of these 3 combinations is equal 12 and I understand that because z= r (Free variable) What's frustrating me it the 8≤ z ≤ 10.

hero (hero):

z = 9, x = 1, y = 2 should be the only true solution.

OpenStudy (anonymous):

i dont see why x/x cant be zero?

OpenStudy (anonymous):

I think i got it! Could it be that because x= -8 + z and y= 20 - 2z We solve by finding z which equals to z=8 and z=10, which means that the answer for z is 8≤ z ≤10! And then we substitute in the given equations to find how many planes we should lease! I think I got it :D

OpenStudy (anonymous):

I just realized I said I think I got it twice. Lol, sorry. it's cause i've been stuck on it for a while.

OpenStudy (anonymous):

Thank you so much for your help!